OpenStudy (anonymous):

Fine an equation of the line having the given slope and containing the given point. m = 8/9, (4, - 1)

OpenStudy (unanimoose):

y = mx + b M = slope. n = y-intercept. Can you fill these in?

OpenStudy (unanimoose):

I'm sorry, b is the y-intercept. Not n. :)

OpenStudy (anonymous):

The formula is I use for this is: y - y1 = m(x - x1)

OpenStudy (unanimoose):

In this case: m = slope x1 = x-pont. y1 = y-point Try now. :)

OpenStudy (anonymous):

\[y = \frac{ 8x }{ 9 } - \frac{ 41 }{ 9 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Are there answer choices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Are you asked to find the equation with Point-Slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If so then the answer would be . . . \[y = \frac{ 8x }{ 9 } - \frac{ 41 }{ 9 }\]

OpenStudy (unanimoose):

Yes, that's correct. But next time walk him through it rather than giving him the direct answer. :)